OpenStudy (anonymous):

I NEED HELP D; Find the variance and standard deviation. {3, 3, 4, 5, 5}

OpenStudy (anonymous):

first find the mean

OpenStudy (anonymous):

mean = (sum of observation)/number of observation For this problem, its (3+3+4+5+5)/5 = 20 /5 = 4

OpenStudy (anonymous):

now, variance is equal to ( sum of (observation - mean )^2 ) / (number of observation - 1)

OpenStudy (anonymous):

variance = 1+1+0+1+1 / 4 = 1

OpenStudy (anonymous):

deviation = average size of deviation = sqrt 1